# Learnable quantum spectral filters for hybrid graph neural networks

### v1

@@ 1.1 Motivation
The need for more computational resources and computational power of quantum computers. With the advancement in machine learning and AI technologies, the number of new papers in computer science increases almost exponentially each year (see arXiv new submission statistics [1]). One of the main triggers for this trending appeal is the recent AI tools based on large language models such as GPT [2], DeepSeek [3], and others [4]. These models are trained using billions of parameters and require huge computational power. Therefore, we can view these models as representing solution space by using large numbers of parameters and thus requiring huge computational power that can be only sustained by a few big companies and wastes energy resources [5, 6]. On the other hand, it is well known that many difficult problems can be approximated by using randomization (see e.g. [7, 8, 9]) and approximation (see e.g. [10, 11]) techniques. While the solution space for many difficult problems grows exponentially [12], the classical computational power of a classical computer increases linearly with the number of CPUs (see Amdahl’s law [13, 14]). Therefore, sustainable improvement and scalability in these models either will either be based on simplification as done by DeepSeek-R1 [3], QWen3 [15], and other similar models or rely on using technologies such as quantum computers [16] that can increase their computational space exponentially in the number of qubits: For instance, while a two-qubit quantum computer deals with a $4\times 4$ matrix, a three-qubit quantum computer uses an $8\times 8$ matrix. This motivates further research in designing models based on quantum computing even if they do not provide a theoretical computational advantage over the classical computers and their expressive powers [17, 18, 19] can be simulated on classical computers [20].
@@ 2 Background and related works
For a given weight matrix $W$ for a graph $G=(V,E)$ with the set of vertices $V$ and edges $E$, the unnormalized Laplacian is defined as $L=D-W$ while the normalized symmetric version is given bby $L=I-D^{-1/2}WD^{-1/2}$. Here, $D$ is the diagonal matrix with elements defined as $d_{ii}=\sum^{N}_{j=1}w_{ij}$ where $w_{ij}$ are the elements of $W$.The Laplacian operator captures many essential graph properties [55, 56] and find applications in many areas [57] such as the spectral clustering [58] or its quantum version [59].
Geometrically motivated data analysis approaches typically relies on the assumption that high-dimensional data lies on a low-dimensional manifold [60]. In this framework, the graph Laplacian corresponds to the Laplacian differential operator on the manifold [61]. While principal component analysis approximate a high dimensional data by identifying a lower dimensional subspace, nonlinear techniques such as the Laplacian eigenmaps transform the data into a lower-dimensional subspace while preserving the geometric relationships. Thus, many properties of the graph can be understood and manipulated through the eigenspectrum of the Laplacian [19, 61].. For instance, for the graph Laplacian, eigenvectors form a Fourier-like basis, with eigenvalues corresponding to frequencies. Therefore, a low-pass spectral filter can be used to smooth signals by attenuating high-frequency components [62].
@@ 2.1 Related works
In addition, there are quantum feature‐map extraction schemes [96, 97, 98] proposed to be used along with CNNs. A more similar work to our approach is Ref. [99], which uses a variational circuit with amplitude embedding and a convolution scheme based on the adjacency matrix to define hybrid quantum CNNs tested on high-energy physics data. The difference in our approach is that we use the parameterized quantum Fourier transform and qubit connections described for a hybrid graph neural network as explained in the Methods. Note that there is also an earlier work [100] that uses the physical realization of a quantum particle as a graph and implements a spectral filter on it.
@@ 3.1.1 Limitations of this approach and possible remedies
t is possible to face issues when the graph is too sparse, so that the adjacency matrix has a row or column of zeros (isolated nodes). In addition, this structure leads to an uninformative qubit connection matrix: for instance, when there are only 1s in the first row/column of the matrix, the phase connection is 0. This can cause the phase matrix (used for $CRZ$ gates) to be zero, rendering the parameterized QFT ineffective.

### v2

@@ 1.1 Motivation
The need for more computational resources and computational power of quantum computers. With the advancement in machine learning and AI technologies, the number of new papers in computer science grows almost exponentially each year (see arXiv new submission statistics [1]). One of the main triggers for this trending appeal is the recent AI tools based on large language models such as GPT [2], DeepSeek [3], and others [4]. These models are trained on billions of parameters and demand enormous computational power. These models can be viewed as representing solution space by using large numbers of parameters. This requires computational resources that only a few big companies sustain and results in substantial waste of energy resources [5, 6]. On the other hand, it is well known that many difficult problems can be approximated by using randomization (see e.g. [7, 8, 9]) and approximation (see e.g. [10, 11]) techniques. While the solution space for many difficult problems grows exponentially [12], the classical computational power of a classical computer increases linearly with the number of CPUs (see Amdahl’s law [13, 14]). Therefore, sustainable improvement and scalability in these models will either be based on simplification as done by DeepSeek-R1 [3], QWen3 [15], and other similar models or rely on using technologies such as quantum computers [16] that can increase their computational space exponentially in the number of qubits: For instance, while a two-qubit quantum computer deals with a $4\times 4$ matrix, a three-qubit quantum computer uses an $8\times 8$ matrix. This motivates further research in designing models based on quantum computing even if they do not provide a theoretical computational advantage over the classical computers and their expressive powers [17, 18, 19] can be simulated on classical computers [20].
@@ 2 Background and related works
For a given weight matrix $W$ for a graph $G=(V,E)$ with the set of vertices $V$ and edges $E$, the unnormalized Laplacian is defined as $L=D-W$ while the normalized symmetric version is given by $L=I-D^{-1/2}WD^{-1/2}$. Here, $D$ is the diagonal matrix with elements defined as $d_{ii}=\sum^{N}_{j=1}w_{ij}$ where $w_{ij}$ are the elements of $W$.The Laplacian operator captures many essential graph properties [55, 56] and find applications in many areas [57] such as the spectral clustering [58] or its quantum version [59].
Geometrically motivated data analysis approaches typically relies on the assumption that high-dimensional data lies on a low-dimensional manifold [60]. In this framework, the graph Laplacian corresponds to the Laplacian differential operator on the manifold [61]. While principal component analysis approximate a high dimensional data by identifying a lower dimensional subspace, nonlinear techniques such as the Laplacian eigenmaps transform the data into a lower-dimensional subspace while preserving the geometric relationships. Thus, many properties of the graph can be understood and manipulated through the eigenspectrum of the Laplacian [19, 61]. For instance, for the graph Laplacian, eigenvectors form a Fourier-like basis, with eigenvalues corresponding to frequencies. Therefore, a low-pass spectral filter can be used to smooth signals by attenuating high-frequency components [62].
@@ 2.1 Related works
In addition, there are quantum feature‐map extraction schemes [96, 97, 98] proposed to be used along with CNNs. A more similar work to our approach is Ref. [99], which uses a variational circuit with amplitude embedding and a convolution scheme based on the adjacency matrix to define hybrid quantum CNNs tested on high-energy physics data. The difference in our approach is that we use the parameterized quantum Fourier transform and qubit connections described for a hybrid graph neural network as explained in the Methods. Note that there is also an earlier work [100] that uses the physical realization of a quantum particle as a graph and implements a spectral filter on it and there are recent studies using quantum Fourier based circuits for the quantum simulation of antisymmetric matrices [101] and for finding partial differential equations in Fourier space [102].
@@ 3.1.1 Limitations of this approach and possible remedies
It is possible to face issues when the graph is too sparse, so that the adjacency matrix has a row or column of zeros (isolated nodes). In addition, this structure leads to an uninformative qubit connection matrix: for instance, when there are only 1s in the first row/column of the matrix, the phase connection is 0. This can cause the phase matrix (used for $CRZ$ gates) to be zero, rendering the parameterized QFT ineffective.